\begin{thebibliography}{ABC}
	
\bibitem{AKS} { E. Ardonne, R. Kedem, M. Stone}, \textit{Fermionic characters and arbitrary highest-weight integrable $\gsl_{r+1}$-modules},
Comm. Math. Phys. {\bf 264} (2006), 427--464.

\bibitem{Ba} I. Baranovi\'c,  \textit{Combinatorial bases of Feigin-Stoyanovsky's type subspaces of level $2$ standard modules for $D_4^{(1)}$}, 
Comm. Algebra {\bf 39} (2011), 1007--1051.

	
	\bibitem{BPT} I. Baranovi\' c, M. Primc, G. Trup\v cevi\' c, \textit{Bases of Feigin–Stoyanovsky’s type subspaces for $C_\ell^{(1)}$}, Ramanujan J (2016). doi:10.1007/s11139-016-9840-y 1007-1051.
	
\bibitem{Bu1} M. Butorac, \textit{Combinatorial bases of principal subspaces for the affine Lie algebra of type $B_2^{(1)}$},
J. Pure Appl. Algebra, {\bf 218} (2014), 424--447.

\bibitem{Bu2} M. Butorac, \textit{Marijana Butorac, Quasi particle bases of principal subspaces of the affine Lie algebra of type $G_2^{(1)}$}, 
Glas. Mat. {\bf  52} (2017), 81--100.
	
\bibitem{C1} { C. Calinescu}, \textit{Intertwining vertex operators and certain representations of $\widehat{\mathfrak{sl}(n)}$}, Commun. Contemp. Math. {\bf 10} (2008), 47--79.


\bibitem{CalLM1} { C. Calinescu, J. Lepowsky, A. Milas}, \textit{Vertex-algebraic structure of the principal subspaces of certain $A_1^{(1)}$-modules, I: level one case}, Int. J. Math. {\bf 19} (2008), 71--92.

\bibitem{CalLM2} { C. Calinescu, J. Lepowsky, A. Milas}, \textit{Vertex-algebraic structure of the principal subspaces of certain $A_1^{(1)}$-modules, II: higher-level case}, J. Pure Appl. Algebra, {\bf 212} (2008), 1928--1950.

\bibitem{CalLM3} { C. Calinescu, J. Lepowsky, A. Milas}, \textit{Vertex-algebraic structure of the principal subspaces of level one modules for the untwisted affine Lie algebras of types A,D,E}, J. Algebra {\bf 323} (2010), no. 1, 167--192.

\bibitem {CLM1}    { S. Capparelli, J. Lepowsky, A. Milas},
\textit{The Rogers-Ramanujan recursion and intertwining operators},
Comm. Contemporary Math. {\bf 5} (2003), 947--966.

\bibitem {CLM2}    { S. Capparelli, J. Lepowsky, A. Milas},
\textit{The Rogers-Selberg recursions, the Gordon-Andrews identities and intertwining operators},  Ramanujan J.  {\bf 12}  (2006),  no. 3, 379--397.

	\bibitem{FS} A. V. Stoyanovsky and B. L. Feigin, \textit{Functional models of the representations of current algebras, and semi-infinite Schubert cells}, (Russian) Funktsional. Anal. i Prilozhen. {\bf 28} (1994), 68--90; translation in Funct. Anal. Appl. {\bf 28} (1994), 55-72; preprint B. Feigin and A. Stoyanovsky, \textit{Quasi-particles models for the representations of Lie algebras and geometry of flag manifold}, hep-th/9308079, RIMS 942.

\bibitem{FJLMM} { B. Feigin, M. Jimbo, S. Loktev, T. Miwa, E. Mukhin}, \textit{Bosonic formulas for $(k,\ell)$-admissible partitions},  Ramanujan J. {\bf 7}  (2003),  no. 4, 485--517.; \textit{Addendum to `Bosonic formulas for $(k,\ell)$-admissible partitions'},  Ramanujan J. {\bf 7}  (2003),  no. 4, 519--530.

\bibitem{FJMMT} { B. Feigin, M. Jimbo, T. Miwa, E. Mukhin, Y. Takeyama}, \textit{Fermionic formulas for $(k,3)$-admissible configurations}, Publ. RIMS {\bf 40} (2004), 125--162.

\bibitem {G} { G. Georgiev},
\textit{Combinatorial constructions of modules for
	infinite-dimensional Lie algebras, I. Principal subspace},  J. Pure
Appl. Algebra \textbf{112} (1996), 247--286.

	
	\bibitem{H} J. Humphreys
	\textit{Introduction to Lie Algebras and Representation Theory}, Springer, New-York, 1994.
	
\bibitem{J1} M. Jerkovi\'c,
\textit{Recurrence relations for characters of affine
	Lie algebra $A_{\ell}^{(1)}$}, J. Pure Appl. Algebra {\bf 213},
913--926.

\bibitem{JP} { M. Jerkovi\'c, M. Primc}, \textit{Quasi-particle fermionic formulas for $(k, 3)$-admissible configurations}, 
Cent. Eur. J. Math. {\bf 10} (2011), 703--721.

	\bibitem{K} {V. G. Kac},
	\textit{Infinite-dimensional Lie algebras}, 3rd ed. Cambridge
	University Press, Cambridge, 1990.
%	
\bibitem{LL} J. Lepowsky, H.-S. Li,
\textit{ Introduction to Vertex Operator Algebras
	and Their Representations}, Progress in Math. {\bf 227},
Birkh\"auser, Boston, 2004

	\bibitem{LP}
	J. Lepowsky and M. Primc,
	\textit{Structure of the standard modules for
		the affine Lie algebra $A_{1}^{(1)}$}, {\em Contemporary Math.}
	{\bf 46}, (1985), 1--84.
%	
%	
\bibitem{MP} A. Meurman and M. Primc, \textit{Annihilating fields of standard modules of $\mathfrak{sl}(2,\mathbb{C})^{\widetilde{} }$ and combinatorial identities}, Memoirs Amer. Math. Soc. {\bf 652}, 1999.

\bibitem{Pe} M. Penn, \textit{Lattice Vertex Superalgebras, I: Presentation of the Principal Subalgebra},
Comm.Alg. {\bf 42} (2014), 933--961.


	\bibitem{P1} M. Primc, \textit{ Vertex operator construction of standard modules
		for $A_n^{(1)}$}, Pacific J. Math {\bf 162} (1994), 143--187.
	
	\bibitem{P2} M. Primc, \textit{ Basic Representations for classical
		affine Lie algebras}, J. Algebra {\bf 228} (2000), 1--50.
	
	\bibitem{P3} M. Primc, \textit{ Combinatorial bases of modules
		for affine Lie algebra $B_2^{(1)}$}, Cent. Eur. J. Math. {\bf 11} (2013), 197--225. 
%	
%%	\bibitem[P\v S]{PS} M. Primc and T. \v Siki\' c, \textit{Leading terms of relations for standard modules of affine Lie algebras $C_{n}\sp{(1)}$},
%%	arXiv:1506.05026 [math.QA]
	
\bibitem{S1} { C. Sadowski}, \textit{Presentations of the principal subspaces of the higher level $\widehat{\gsl (3)}$-modules}, J. Pure
Appl. Algebra, {\bf 219} (2015) 2300--2345.

\bibitem{Sto} A.V. Stoyanovsky, \textit{Lie Algebra Deformations and Character Formulas}
(Russian) Funktsional. Anal. i Prilozhen. {\bf 32} (1998), 84--86; translation in Funct. Anal. Appl. {\bf 32} (1998),  66--68.



\bibitem{T1} G. Trup\v cevi\' c, \textit{Combinatorial bases of Feigin-Stoyanovsky's type subspaces of level $1$ standard $\tilde{\gsl}(\ell+1,\C)$-modules}, Comm. Algebra {\bf 38} (2010),
3913--3940.

\bibitem{T2} G. Trup\v cevi\' c, \textit{Bases of standard modules for affine Lie algebras of type $C_\ell^{(1)}$}, preprint,	arXiv:1706.10089 [math.QA].

\end{thebibliography}

\documentclass{amsart}
\usepackage[croatian]{babel}

\usepackage{amsmath, amssymb, amsfonts}
\usepackage[all]{xy}
\usepackage{enumitem}   
\usepackage{float}

\newtheorem{tm}{Theorem}

\newenvironment{dokaz}
{\noindent\emph{Proof:}\ }
{\hfill $\blacksquare$}

\newcommand{\Z}
{{\mathbb Z}}
\newcommand{\N}
{{\mathbb N}}
\newcommand{\C}
{{\mathbb C}}
\newcommand{\m}
{{\mathbf m}}

\newcommand{\g}
{{\mathfrak g}}

\newcommand{\gt}
{\tilde{{\mathfrak g}}}

\newcommand{\h}
{{\mathfrak h}}
\newcommand{\n}
{{\mathfrak n}}
\newcommand{\gsl}
{{\mathfrak sl}}

\newcommand{\Gamt}
{\tilde{{\Gamma}}}

\begin{document}

\title[Presentations of Feigin-Stoyanovsky's type subspaces]{Presentations of Feigin-Stoyanovsky's type subspaces of standard modules for affine Lie algebras of type $C_\ell^{(1)}$}
%\Country{University of Zagreb, Croatia}
\author{Goran Trup\v{c}evi\'{c}}
\address{Faculty of Teacher Education, University of Zagreb, Zagreb, Croatia}
%%\curraddr{}
\email{goran.trupcevic@ufzg.hr}
\thanks{Partially supported by the Croatian Science Foundation grant 2634.}
\subjclass[2000]{Primary 17B67; Secondary 17B69, 05A19.}
\keywords{affine Lie algebras, principal subspaces, generators and relations}
%%\date{}
%%\dedicatory{}
%
\begin{abstract}
	Feigin-Stoyanovsky's type subspace $W(\Lambda)$ of a standard $\gt$-module $L(\Lambda)$ is a $\gt_1$-submodule of $L(\Lambda)$ generated by
	 the highest-weight vector $v_\Lambda$, 	where $\gt_1$ is a certain commutative subalgebra of $\gt$. Based on the description of basis of 
	 $W(\Lambda)$ for $\gt$ of type $C_\ell^{(1)}$, we 	give a presentation of this subspace in terms of generators and relations
	$$W(\Lambda)\simeq U(\gt_1^-)/J.$$
\end{abstract}

\maketitle

\section{Introduction}

B. Feigin and A. Stoyanovsky introduced principal subspaces of standard modules for affine Lie algebras of type $A_1^{(1)}$ and $A_2^{(1)}$ 
in \cite{FS} where they have recovered Rogers-Ramanujan type identities by considering graded dimensions of these subspaces. An important part
of their investigation was the knowledge of presentations of these subspaces in terms of generators and relations. 
Another type of principal subspaces, called  Feigin-Stoyanovsky's type subspace, was introduced by M. Primc who constructed bases of these subspaces
in different cases (\cite{P1}, \cite{P2}, \cite{P3}, \cite{JP}). 
These kind of subspaces were further studied by many authors (\cite{Sto}, \cite{G}, \cite{CLM1}, \cite{CLM2}, \cite{AKS}, \cite{FJLMM}, \cite{FJMMT}, \cite{Ba}, \cite{Bu1}, \cite{Bu2}, \cite{J1}, \cite{T1}, \dots) and the knowledge of presentation
presents an important question in this study (\cite{C1},  \cite{CalLM1}, \cite{CalLM2}, \cite{CalLM3}, \cite{S1}, \cite{Pe}).

In our previous works  we have described bases of Feigin-Stoyanovsky's type subspaces  of standard modules for affine Lie algebras 
of type $C_\ell^{(1)}$ (\cite{BPT}) and obtained from them basis for the whole standard modules (\cite{BPT}). In this note we use the description of bases of a Feigin-Stoyanovsky's type subspaces to give presentations of these subspaces in terms of generators and relations.

\section{Feigin-Stoyanovsky's type subspaces}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%% prosta kondim Liejeva algebra

Let ${\mathfrak g}$ be a complex simple  Lie algebra of type $C_\ell$ with a Cartan
subalgebra $\h$ and a root decomposition ${\mathfrak g}={\mathfrak h}+\sum {\mathfrak g}_\alpha$. Let 
$$
R=\{\pm \epsilon_i\pm \epsilon_j\,|\, 1\leq i\leq j\leq \ell\}\backslash\{0\}
$$ be the corresponding root system realized in $\mathbb R\sp\ell$ with
the canonical basis $\epsilon_1,\dots,\epsilon_\ell$.
Fix simple roots
$$\alpha_1=\epsilon_1-\epsilon_2,\quad\dots,\quad\alpha_{\ell-1}=\epsilon_{\ell-1}-\epsilon_\ell,\quad\alpha_\ell=2\epsilon_\ell$$ 
and let   $ \g=\n_-+ \h + \n_+$ be the corresponding triangular decomposition. Let
$\theta=2\alpha_1+\dots+2\alpha_{\ell-1}+\alpha_\ell=2\epsilon_1$ be the maximal root and
$$
\omega_r= \epsilon_1+\dots+ \epsilon_r, \qquad r=1,\dots,\ell
$$ 
fundamental weights (cf. \cite{H}). 
Fix root vectors $x_\alpha\in\g_\alpha$. % and denote by $\alpha\sp\vee\in\mathfrak h$ dual roots. 
We identify $\mathfrak h$ and $\mathfrak h\sp*$ via the Killing form $\langle\,,\,\rangle$ normalized in 
such a way that  $\langle\theta,\theta\rangle=2$.

\bigskip

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%% Afina Liejeva algebra

Let $\tilde{\mathfrak g}$ be the affine Lie algebra of type $C_\ell\sp{(1)}$ associated  to $\g$,
$$
\gt = \mathfrak{g} \otimes
\mathbb{C}[t,t^{-1}] + \mathbb{C}c, + \mathbb{C} d,
$$
with the canonical central element $c$ and the degree element $d$ (cf. \cite{K}). Let  
$$
\gt=\tilde{\n}_-+ \tilde{\h} + \tilde{\n}_+,
$$
be a triangular decomposition of $\gt$, where $
\tilde{\n}_-=\n_-+ \g\otimes t^{-1}\C [t^{-1}]$, $\tilde{\h}=\h  + \mathbb{C}c + \mathbb{C} d$, $\tilde{\n}_+=\n_++ \g\otimes t\C [t]$.
Denote by $\Lambda_0,\dots,\Lambda_\ell$ fundamental weights of $\gt$.
 
For $x\in{\mathfrak g}$ and $n\in\mathbb Z$ denote by $x(n)=x\otimes t^{n}$ and $x(z)=\sum_{n\in\mathbb Z}
x(n) z^{-n-1}$, where $z$ is a  formal variable. 

Let $L(\Lambda)$ be a standard $\tilde{\mathfrak g}$-module with the highest weight
$$
\Lambda=k_0 \Lambda_0+k_1 \Lambda_1+\dots+k_\ell \Lambda_\ell,
$$
$k_i\in\Z_+$ for $i=0,\dots,\ell$, and fix a highest weight vector $v_\Lambda$.
Denote by
$k=\Lambda(c)$ the level of  $\tilde{\mathfrak g}$-module $L(\Lambda)$,
$
k=k_0 +k_1 +\dots+k_\ell$.

\bigskip

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%% Z-gradacija

Fix the minuscule weight 
$\omega=\omega_\ell=\epsilon_1+\dots+\epsilon_\ell \in\h^*$; then $\langle\omega,\alpha\rangle\in\{-1,0,1\}$ for all $\alpha\in R$ and
define {\it the set of colors} 
$$\Gamma  =
\{\,\alpha \in R \mid \langle\omega,\alpha\rangle = 1\}  =  \{\epsilon_i+\epsilon_j\,|\, 1\leq i \leq j\leq \ell\}.
$$
Write
$$(ij)=\epsilon_i+\epsilon_j\in\Gamma\quad\textrm{and}\quad x_{ij}=x_{\epsilon_i+\epsilon_j}.$$

This gives a $\mathbb Z$-gradation of $\gt$; let ${\mathfrak g}_0 = {\mathfrak h} +
\sum_{\langle\omega,\alpha\rangle=0}\, {\mathfrak g}_\alpha$, then
$$
\gt=\gt_{- 1}+\gt_{0}+\gt_{ 1},
$$
where
$$
\gt_0 = \g_0 \otimes\C [t,t^{-1}] \oplus \C c \oplus \C d,\qquad
\gt_{\pm 1} = \sum_{\alpha \in \pm\Gamma}\, {\mathfrak g}_\alpha \otimes\C [t,t^{-1}].
$$
The subalgebra $\gt_1$ is commutative, and $\g_0$ acts on $\gt_1$ by adjoint action.

\bigskip

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%% Feigin-Stoyanovsky subspace

{\it Feigin-Stoyanovsky's type subspace} of $L(\Lambda)$ is a $\gt_1$-submodule of $L(\Lambda)$ generated by the highest-weight vector $v_\Lambda$, 
$$
W(\Lambda) = U(\gt_1)\cdot v_\Lambda = U(\gt_1^-)\cdot v_\Lambda\subset L(\Lambda),
$$
where $\gt_1^-=\gt_1 \cap \tilde{\n}_-$.

\bigskip

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%% DC i IC

A monomial $\m\in U(\gt_1^-)= S(\gt_1^-)$,
$$\m=\cdots x_{i'
	j'}(-n-1)^{b_{i'j'}}\cdots x_{i j}(-n)^{a_{ij}}  \cdots,$$% \quad\in U(\gt_1^-)$$ 
is said to satisfy {\em 	difference conditions} for $W(\Lambda)$, $DC$ for short,  if 
for any $n\in\N$ and $i_1 \leq \dots \leq i_t \leq j_t\leq\ldots\leq j_1 \leq i_{t+1} \leq \dots \leq i_{s} \leq j_{s} \leq \ldots\leq j_{t+1}$,
$$ %\label{DC_def}
b_{i_1 j_1}+\ldots+b_{i_t j_t}+a_{i_{t+1} j_{t+1}}+\ldots+a_{i_s
	j_s}\leq k,
$$
where $a_{ij}$'s and $b_{ij}$'s denote exponents of $x_{ij}(-n)$ and
$x_{ij}(-n-1)$ in $\m$, respectively.

A monomial $\m$ satisfies {\em initial conditions} for $W(\Lambda)$, $IC$ for short, if 
for every $i_1 \leq \dots \leq i_t \leq j_t\leq\ldots\leq j_1$,
$$ %\label{IC_def}
a_{i_1 j_1}+\ldots+a_{i_t 	j_t}\leq k_0+k_1+\ldots + k_{j_1-1}
$$
where $a_{ij}$'s denote exponents of $x_{ij}(-1)$ in $\m$.

\begin{tm}[{\cite{BPT}}]%\label{FSbaza_tm}
	The set
	$$\{\m v_\Lambda \,|\, \m \ \textrm{satisfies}\ DC\ \textrm{and}\ IC\ \textrm{for}\ W(\Lambda)\}$$
	is a basis of $W(\Lambda)$.
\end{tm}

\section{Presentation of Feigin-Stoyanovsky's type subspaces}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%% Relacije DC

Difference conditions are consequences of the adjoint action of $\g_0$ on the vertex-operator relation
$$x_\theta(z)^{k+1}=0,$$ 
or, equivalently, on a family of relations
$$\sum_{\substack{n_1,\dots,n_{k+1}\geq 1 \\ n_1+\dots+n_{k+1}=N}}x_{11}(-n_1)\cdots x_{11}(-n_{k+1})=0,\quad \textrm{for}\ N\geq k+1$$
on $L(\Lambda)$ (cf. \cite{BPT}; see also  \cite{LP}, \cite{MP}, \cite{LL})). 

Root vectors of $\g$ can be chosen so that the action of $\g_0$ on $\g_1$ is given by 
$$
 [x_{-\alpha_i},x_{ij}]=x_{i+1,j}, \quad [x_{-\alpha_j},x_{ij}]=x_{i,j+1},\quad [x_{-\alpha_i},x_{ii}]= 2x_{i,i+1}
$$
(cf. \cite{H}). Then one easily sees that the adjoint action gives the following family of relations on $L(\Lambda)$:
$$
%\label{DC_rel}
\sum_{\substack{n_1+\dots+n_{k+1}=N\\ \{i_1,\dots,i_{k+1},j_1,\dots,j_{k+1} \}=\{1^{m_1},\dots,\ell^{m_\ell}\} }} 
\hspace{-3ex} C_{\bf ij} x_{i_1 j_1}(-n_1)\cdots x_{i_{k+1} j_{k+1}}(-n_{k+1})=0,
$$
for some nonnegative integers $C_{\bf ij}$, where the sum runs over all such partitions ${\bf i},{\bf j}$ of a multiset $\{1^{m_1},\dots,\ell^{m_\ell}\}$, $m_1+\dots+m_\ell=2(k+1)$ (cf. \cite{BPT}).
 
One obtains the difference conditions by finding minimal monomials of these relations,
the so called {\em leading terms} of relations, whose multiples can be excluded from the spanning set 
For this, a linear order on monomials is introduced.
Define a linear order on the set of colors $\Gamma$:
$(i'j')<(ij)$ if $i'>i$ or $ i'=i,\, j'>j$. 
On the {\em set of variables} $\Gamt^-=\{x_\gamma(n) \,|\,\gamma\in\Gamma, n\in \Z_- \}$  define a linear order by
$x_\alpha(n)<x_\beta(n')$ if $n<n'$ or
$n=n',\, \alpha<\beta$. 
For monomials, assume that factors descend from right
to left, then use a lexicographic order (compare factors the greatest to the lowest one).
Order $<$ is compatible with multiplication (see \cite{P1}, \cite{T1}): if $\m_1<\m_2$ then $\m \m_1<\m \m_2$, for  $\m,\m_1, \m_2 \in U(\gt_1^-)$.

\bigskip

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%% Relacije IC

For initial conditions consider decompositions 
%$$\Lambda=\Lambda^{(r)}+\Lambda_{(r)}, \qquad \Lambda^{(r)}=\sum_{t< r } k_t \Lambda_t,\quad \Lambda_{(r)}=\sum_{t\geq r}  k_t \Lambda_t.$$
$$\Lambda=\Lambda^{(r)}+\Lambda_{(r)}, \qquad \Lambda^{(r)}=k_0 \Lambda_0+\dots+k_{r-1}\Lambda_{r-1},\quad \Lambda_{(r)}=k_r \Lambda_r+\dots + k_\ell \Lambda_\ell,$$
for $1\leq r \leq \ell$. 
By $v^{(r)}$ and $v_{(r)}$ denote highest weight vectors of the 
associated standard modules $L(\Lambda^{(r)})$ and $L(\Lambda_{(r)})$ of level
$k^{(r)}=k_0+\dots+k_{r-1} $ and $k_{(r)}=k_{r}+\dots+k_\ell$, respectively. Then $L(\Lambda)$ can be embedded in a tensor product 
$L(\Lambda)\subset L(\Lambda^{(r)}) \otimes L(\Lambda_{(r)})$.  
Since $x_{ij}(-1) v_{\Lambda_r} = 0$ if and only if $j\leq r$ (cf. \cite{BPT}), we have
$$\m (v^{(r)}\otimes v_{(r)}) = (\m v^{(r)}) \otimes v_{(r)}$$
for $\m=x_{i_1 j_1}(-1) \cdots x_{i_t j_t}(-1)$ such that $j_s \leq r$, $1\leq s \leq t$.
Hence, relations between such monomials corresponding to difference conditions for $W(\Lambda^{(r)})$ automatically become 
relations in $L(\Lambda)$ (cf. \cite{BPT}).
This gives the following family of relations on $L(\Lambda)$: 
\begin{equation}
\label{IC_rel}
\sum_{\{i_1,\dots,i_{k+1},j_1,\dots,j_{k+1} \}=\{1^{m_1},\dots,r^{m_r}\} } 
\hspace{-3ex} C_{\bf ij} x_{i_1 j_1}(-1)x_{i_2 j_2}(-1)\cdots x_{i_{k^{(r)}+1} j_{k^{(r)}+1}}(-1)=0,
\end{equation}
for some nonnegative integers $C_{\bf ij}$, where the sum runs over all such partitions ${\bf i}, {\bf j}$ of a multiset $\{1^{m_1},\dots,r^{m_r}\}$, $m_1+\dots+m_r=2(k^{(r)}+1)$.

Alternatively, for $r\geq 2$ let $\g_{(r)}\subset \g_0$ be the subalgebra generated by elements $x_{\pm\alpha_t}$, $1\leq t < r$.
Start from a relation 
$$x_{11}(-1)^{k^{(r)}+1} (v^{(r)}\otimes v_{(r)}) = (x_{11}(-1)^{k^{(r)}+1} v^{(r)})\otimes v_{(r)} = 0.$$
Now the adjoint action of  $\g_{(r)}$ on the above relation gives  relations \eqref{IC_rel}. For $r=1$,  relations \eqref{IC_rel} come down to only one relation
$$x_{11}(-1)^{(k_0+1)} (v^{(1)}\otimes v_{(1)})=0.$$

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%% Prezentacija

\bigskip

Recall that Feigin-Stoyanovsky's type subspace $W(\Lambda)$ is 
$$W(\Lambda)=U(\gt_1^-)\cdot v_\Lambda.$$
Since $\gt_1$ is commutative, universal enveloping algebra of $\gt_1^-$ is isomorphic to a polynomial algebra $\C[\Gamt^-]$. Hence, there is a 
surjection $$f_\Lambda:\C[\Gamt^-]\to W(\Lambda), \qquad f:\m \to \m\cdot v_\Lambda.$$
We want to describe the kernel of this map, $\ker f_\Lambda\subset \C[\Gamt^-]$, so that
$$W(\Lambda) \simeq \C[\Gamt^-]/\ker f_\Lambda, $$
as vector spaces.

\begin{tm}
Let $J_\Lambda\subset\C[\Gamt^-]$ be the ideal generated by the following sets
\begin{eqnarray*} 
%\label{Pres_eq1} 
U(\g_0)\cdot\left(\sum_{\substack{n_1,\dots,n_{k+1}\geq 1 \\ n_1+\dots+n_{k+1}=N}}x_{11}(-n_1)\cdots x_{11}(-n_{k+1})\right),& &  \textrm{for}\ N\geq k+1,\\
%\label{Pres_eq2} 
U(\g_{(r)})\cdot x_{11}(-1)^{k^{(r)}+1},& &  \textrm{for}\ r=2,\dots,\ell,\\
%\label{Pres_eq3} 
x_{11}(-1)^{k_0+1}.& &   
\end{eqnarray*}
Then $\ker f_\Lambda = J_\Lambda$.
\end{tm}

\begin{dokaz}
The generators of $J_\Lambda$ are exactly the elements that appear on the left side of relations on $W(\Lambda)$. Therefore they
lie in the kernel of $f_\Lambda$, and we can factorize $f_\Lambda$
to a quotient map
$$	\bar{f_\Lambda}: \C[\Gamt^-]/J_\Lambda \to  W(\Lambda).	$$ 
This map is clearly a surjection, since $f_\Lambda$ is a surjection. 

We can imitate the proof for the spanning set for $W(\Lambda)$ (cf. Proposition 2 and 4 in \cite{BPT}) to reduce the spanning set for $\C[\Gamt^-]/J_\Lambda$. 
Consider the generators of $J_\Lambda$ and identify the minimal monomial inside each one; their multiples can be excluded from the spanning set. 
Like in \cite{BPT},  we get 
	$$\mathcal{B}=\{\m\,|\,\m\ \textrm{satisfies DC and IC for}\ W(\Lambda)\}$$
as a spanning set of $\C[\Gamt^-]/J_\Lambda$.
	
To see that $\bar{f_\Lambda}$ is an injection, note that $\bar{f_\Lambda}$ maps $\mathcal{B}$ bijectively onto 
$$\{\m v_\Lambda\,|\,\m \ \textrm{satisfies DC and IC for}\ W(\Lambda)\}\subset W(\Lambda),$$
which is a basis of $W(\Lambda)$.  This means that $\mathcal{B}$ is also linearly
independent. Hence $\bar{f_\Lambda}$ maps a basis of $\C[\Gamt^-]/J_\Lambda$ onto a basis 
of $W(\Lambda)$ and therefore $\bar{f_\Lambda}$ is a bijection.
\end{dokaz}

\end{document}